What is the capacitance of a parallel plate capacitor with metal plates, each of area 1.00 m2, separated by 1.00 mm?

1
Expert's answer
2022-04-17T17:07:09-0400

The capacitance is given as follows:


"C = \dfrac{\epsilon_0A}{d}"

where "A = 1.00m^2" is the area of the plate, "d = 1.00mm = 1.00\times 10^{-3}m" is the distance between plates, and "\epsilon_0 \approx 8.85\times 10^{-12} F/m" is vacuum permittivity. Thus, obtain:


"C = \dfrac{8.854\times 10^{-12} F/m\cdot 1.00m^2}{1.00\times 10^{-3}m} =8.
85\times 10^{-9} F"

Answer. "8.
85\times 10^{-9} F".